Determine for which value (s) of λ the real part of z =1+λi /1- λi equals zero
1
Expert's answer
2021-07-21T06:58:46-0400

"z=\frac{1+\lambda i}{1-\lambda i} =\frac{(1+\lambda i)^2}{(1-\lambda i)(1+\lambda i)}=\frac{1+2\lambda i-\lambda ^2
}{1+\lambda ^2
}= \frac{1-\lambda^2}{1+\lambda ^2}+\frac{2\lambda }{1+\lambda^2}i"

If "\lambda" is a real number, then "\text{Re} \ z=\frac{1-\lambda^2}{1+\lambda^2}=0"


"1-\lambda^2=0"

"\lambda^2=1"

"\lambda=\pm 1"


Answer: "\lambda =\pm1."